# Benchmarking Pretrained Molecular Embedding Models For Molecular Representation Learning

### v3

@@ Abstract
Pretrained neural networks have attracted significant interest in chemistry and small molecule drug design. Embeddings from these models are widely used for molecular property prediction, virtual screening, and small data learning in molecular chemistry. This study presents the most extensive comparison of such models to date, evaluating 25 models across 25 datasets. Under a fair comparison framework, we assess models spanning various modalities, architectures, and pretraining strategies. Using a dedicated hierarchical Bayesian statistical testing model, we arrive at a surprising result: nearly all neural models show negligible or no improvement over the baseline ECFP molecular fingerprint. Only the CLAMP model, which is also based on molecular fingerprints, performs statistically significantly better than the alternatives. These findings raise concerns about the evaluation rigor in existing studies. We discuss potential causes, propose solutions, and offer practical recommendations.
Introduction
The translation of molecular structures into information-rich numerical representations is a cornerstone of modern chemoinformatics and a critical step for harnessing machine learning in drug discovery and materials science. Molecular representation learning has emerged as a powerful paradigm, moving beyond handcrafted descriptors to automatically learn salient features from molecular data. Recently, the field has seen a surge in the development of pretrained molecular embedding models. This trend mirrors the transformative impact of representation learning in other scientific domains, where models are pretrained on vast unlabeled datasets to produce powerful general-purpose embeddings. Notable successes include models based on self-supervised learning (SSL), such as DINO (Caron et al. 2021) and DINOv2 in computer vision, Sentence Transformers and TSDAE (Wang, Reimers, and Gurevych 2021) for dense text representations in NLP, and multimodal architectures such as CLIP and BLIP-2 that connect images and text. Inspired by these advances, molecular models are similarly trained on massive chemical databases to generate universal embeddings.
@@ Introduction
Molecular Embedding Approaches
@@
Molecular fingerprints are feature extraction methods based on identifying small subgraphs within a molecule and detecting their presence or counting their occurrences, yielding binary and count variants, respectively. They can be broadly classified into substructural and hashed types (Adamczyk and Ludynia 2024). Substructural fingerprints detect predefined patterns, such as functional groups or ring systems, that are typically determined by expert chemists. Hashed fingerprints, on the contrary, define general shapes of extracted subgraphs, convert them into numerical identifiers, and hash them using a modulo function into a fixed-length output vector. Common examples include circular neighborhoods in Extended Connectivity FingerPrint (ECFP) (Rogers and Hahn 2010), paths of length 4 in the Topological Torsion (TT) (Nilakantan et al. 1987), and the shortest paths between atom pairs in Atom Pair (AP) fingerprint.
@@
Although not task-adaptive, hashed fingerprints remain widely used in chemoinformatics and molecular machine learning due to their flexibility, computational efficiency, and consistently strong performance. In many cases, they continue to outperform more complex approaches, such as GNNs (Jiang et al. 2021; Baptista et al. 2022; Adamczyk and Czech 2024; Adamczyk, Ludynia, and Czech 2025).
Graph neural networks
@@
Most Graph Neural Network (GNN) architectures follow a message-passing framework (Gilmer et al. 2017). The initial embedding of each atom consists of elementary chemical descriptors such as the element type or charge. In each GNN layer, an atom receives embeddings from its neighbors (messages) and updates its own embedding accordingly, incorporating this information. Most molecular GNNs also include bond features and embeddings in this process. To obtain a whole-molecule embedding, atom embeddings are aggregated using a readout function, such as a channel-wise average or sum. The message-passing mechanism integrates both functional and structural information about the molecule in an end-to-end learning framework.
@@
Graph Isomorphism Network (GIN) (Xu et al. 2019) is one of the most widely used GNN architectures, as it was proven to be as expressive as the Weisfeiler-Lehman isomorphism test in distinguishing non-isomorphic graphs. GIN uses a two-layer multilayer perceptron (MLP) for atom embedding updates during message passing and employs sum readout. All GNN pretraining methods described below rely on a GIN backbone.
@@
Context Prediction (ContextPred) (Hu et al. 2020) was proposed as an SSL pretraining procedure for message-passing GNNs. For each atom, a $K$-hop neighborhood subgraph is defined and encoded as a vector by the GNN that is being pretrained. A corresponding context subgraph, located between $r_{1}$ and $r_{2}$ hops away, is encoded by a separate GNN. The pretraining objective is binary classification with negative sampling, where the model learns to distinguish correct (positive) neighborhood-context pairs from randomly sampled (negative) ones.
@@
GraphMVP (Liu et al. 2022) combines contrastive and generative SSL, aligning molecular 2D and 3D representations, i.e., topological graphs and conformations. These two views are encoded using GIN and SchNet (Schütt et al. 2018), respectively. In the contrastive setup, positive pairs consist of a molecule and its conformers, while the generative objective minimizes the variational autoencoder (VAE) loss between the two representations.
@@
GraphFP (Luong and Singh 2023) introduces graph fragmentation for pretraining, employing both contrastive and predictive SSL. Frequent subgraph mining is used to decompose molecules into fragments. In contrastive learning, fragments and their constituent atoms form positive pairs, while atoms from unrelated fragments serve as negative pairs. The atom-encoding GNN is pretrained to generate atom-level embeddings (which can later be pooled), while a separate GNN encodes entire fragments. In addition, a predictive task is used to classify the presence of fragments, providing a multitask pretraining signal for the atom-encoding GNN.
@@
MolR (Wang et al. 2022) also uses contrastive learning, but incorporates chemical reaction information. Positive samples are constructed from known reactant-product pairs in reaction databases, while negative samples are created by randomly shuffling these pairs. All molecules are encoded using the same GNN being pretrained, based on the TAG architecture (Du et al. 2017).
@@
GEM (Fang et al. 2022) relies on predictive SSL, primarily using 3D molecular conformations. Pretraining tasks include predicting bond lengths, bond angles, and interatomic distances, with parts of the molecule masked as input. Additional auxiliary tasks involve predicting molecular fingerprint bits (MACCS and ECFP) based on the topological graph. Unlike the other architectures described above, which use only topological graphs at inference time, GEM requires 3D conformers as input, making both pretraining and inference computationally expensive due to the need for conformation generation.
Graph transformers
@@
Graph transformer architectures extend self-attention mechanisms to molecular graphs, replacing or augmenting the localized message-passing of GNNs with global attention layers. Atoms are treated as tokens, and edges are encoded as pairwise biases injected into the attention scores, allowing information to propagate between any two atoms within a single layer. Compared to message-passing GNNs, this design more efficiently captures long-range dependencies and enables the incorporation of rich edge or distance features through additive or multiplicative attention biases.
@@
GROVER (Rong et al. 2020) is a hybrid transformer-GNN model in which the attention heads are biased by chemoinformatics-derived edge features. It is pretrained using two self-supervised tasks: multioutput regression (MTR), predicting eight physicochemical descriptors on masked subgraphs, and multioutput classification (MTC), predicting the presence of functional groups (structural motifs). The graph-level embedding is obtained by concatenating a [CLS] token with pooled atomic embeddings.
@@
MAT (Maziarka et al. 2020) introduces distance-aware attention by incorporating adjacency and shortest-path kernels into the query-key dot product. Pretraining combines masked atom and bond prediction, analogous to masked language modeling (MLM) in NLP, with supervised regression on QM9 quantum-chemical properties, encouraging the encoder to learn both topological context and physicochemical properties.
@@
R-MAT (Maziarka et al. 2024) builds on MAT by introducing relative positional encodings derived from graph distances and ring memberships, improving robustness to molecule size and graph sparsity. Inspired by GROVER, it uses subgraph masking as a pretraining objective, making it more challenging than single-atom masking of MAT, and resulting in superior performance.
@@
Uni-Mol (Zhou et al. 2023) and its extended version, Uni-Mol2 (Ji et al. 2024), apply graph transformers to 3D conformers using SE(3)-equivariant attention mechanisms. The encoder is jointly pretrained to reconstruct masked atoms, predict pairwise distances and angles, and distinguish positive from negative conformer–SMILES pairs. These objectives produce embeddings that integrate both molecular topology and 3D geometry.
Text transformers
@@
Molecules serialized as SMILES (Weininger 1988) or SELFIES (Krenn et al. 2020) strings can be processed directly using NLP-inspired architectures. Although pretraining can incorporate chemistry-specific biases, it is not strictly necessary, and the simplicity of NLP-derived pretraining is often a notable advantage of these models. In encoder–decoder architectures, only the encoder is used after pretraining to compute molecular embeddings.
@@
Continous Data-Driven Descriptors (CDDD) (Winter et al. 2019) is an RNN-based architecture. It was trained using a purely text-based translation task, from random SMILES to canonical SMILES, calculated with RDKit (Landrum et al. 2025).
@@
Chemformer (Irwin et al. 2022) is based on BART encoder-decoder transformer (Lewis et al. 2020), pretrained using denoising SSL objective. Given a SMILES string, it is first randomized (e.g., by starting at a different atom) and some tokens are randomly masked; the model is then trained to reconstruct the original SMILES.
@@
MolBERT (Fabian et al. 2020) is an encoder-only transformer based on the original BERT architecture. It is pretrained using two objectives: masked language modeling (MLM), and multitask regression (MTR), in which the model predicts 200 physicochemical molecular properties computed with RDKit. This dual objective is intended to combine contextual learning from SMILES with molecular-level inductive biases.
@@
ChemBERTa (Chithrananda, Grand, and Ramsundar 2020) uses a RoBERTa-style encoder-only architecture. The authors explore two distinct pretraining variants: one based solely on MLM and another based solely on MTR. The MTR-only variant is reported to be simpler to train compared to MolBERT, and achieves better performance than both the MLM variant and MolBERT combined objective.
@@
MoLFormer (Ross et al. 2022) is another encoder-only SMILES transformer. Its authors aim to further simplify pretraining by relying entirely on a large pretraining corpus (derived from PubChem and ZINC) and standard NLP procedures, rather than domain-specific architectural changes. The model is based on BERT, with modifications including linear attention and rotary position embeddings, and is pretrained using MLM.
@@
SimSon (Lee et al. 2025) uses an encoder-only backbone and is pretrained by contrastive SSL. Randomized SMILES strings of the same molecule form positive pairs, while strings from different molecules form negative pairs. Despite using such a simple pretraining procedure and a much smaller dataset, SimSon is shown to be competitive with the aforementioned architectures.
@@
ChemFM (Cai et al. 2024) is built on an encoder-only LLaMA (Touvron et al. 2023) backbone and is pretrained using GPT-style causal language modeling (CLM). It is significantly larger than other models, with 1B and 3B parameter variants. The molecule embedding is derived from the output vector of the end-of-sequence ([EOS]) token.
@@
ChemGPT (Frey et al. 2023) adopts a decoder-only architecture based on GPT-Neo (Black et al. 2021), and is trained on SELFIES strings using CLM. The authors train multiple model variants up to 1.3B parameters, demonstrating that neural scaling laws hold for chemistry-focused text-based transformers as well.
@@
SELFormer (Yüksel et al. 2023) is based on an encoder-only RoBERTa architecture but uses SELFIES, rather than SMILES. Its authors assume that the more regular syntax of SELFIES facilitates learning and improves the capture of relevant molecular patterns. Pretraining uses the MLM objective.
Hybrid models
@@
Mol2Vec (Jaeger, Fulle, and Turk 2018) combines graph-based ECFP4 descriptors with the text-based Word2Vec approach (Mikolov et al. 2013) to learn context-aware token representations. It first computes integer ECFP identifiers for circular subgraphs centered on each atom, using radii 0 and 1. These subgraphs are treated as words in a sequence, with each one forming a token. A Word2Vec skip-gram model is then used to pretrain context-aware vector representations for each token. During inference, all subgraph vectors are summed to obtain a whole-molecule embedding.
@@
COATI (Kaufman et al. 2024) employs a multimodal encoder-decoder architecture, pretrained using CLIP-style (Radford et al. 2021) multimodal contrastive learning. The 3D molecule conformer is encoded using an E(3)-equivariant GNN (Satorras, Hoogeboom, and Welling 2021), and its SMILES string with a RoFormer (Su et al. 2024). The model is pretrained with two objectives: contrastive learning between those two representations (treating embeddings of the same molecule as positive pairs), and autoregressive SMILES reconstruction using a decoder. At inference time, molecular embeddings are obtained from the SMILES encoder.
@@
CLAMP (Seidl et al. 2023) is a multimodal encoder model trained with supervised contrastive learning based on ChEMBL bioactivity assay data. Each molecule is encoded via the concatenation of two molecular fingerprints (ECFP and RDKit) processed through a two-layer MLP, which serves as the final embedding. Assay descriptions are encoded using Latent Dirichlet Allocation (LDA). Positive pairs consist of a bioactive molecule and the description of the assay in which it is active, while negative pairs involve nonbioactive compounds. This design incorporates a bioactivity-specific inductive bias into the fingerprint-based representation.
Methods

### v4

@@ Abstract
Motivation: Pretrained neural networks have attracted significant interest in chemistry and small molecule drug design. Embeddings from these models are widely used for molecular property prediction, virtual screening, and small data learning in molecular chemistry. However, there is no unified, large-scale comparison of embeddings from those models in literature, making their advantages over each other and over classical molecular fingerprints unclear.
Results: This study presents the most extensive comparison of such models to date, evaluating 25 models across 25 datasets. Under a fair comparison framework, we assess models spanning various modalities, architectures, and pretraining strategies. Using a dedicated hierarchical Bayesian statistical testing model, we arrive at a surprising result: nearly all neural models show negligible or no improvement over the baseline ECFP molecular fingerprint. Only the CLAMP model, which is also based on molecular fingerprints, performs statistically significantly better than the alternatives. These findings raise concerns about the evaluation rigor in existing studies. We discuss potential causes, propose solutions, and offer practical recommendations.
Contact: praski@agh.edu.pl, jadamczy@agh.edu.pl
Supplementary information: All code and data are available at:
https://github.com/scikit-fingerprints/benchmarking_molecular_models
keywords:
molecular property prediction, molecular embeddings, machine learning, chemoinformatics, molecular fingerprints
1 Introduction
The translation of molecular structures into information-rich numerical representations is a cornerstone of modern chemoinformatics and a critical step for harnessing machine learning in drug discovery and materials science. Molecular representation learning has emerged as a powerful paradigm, moving beyond handcrafted descriptors to automatically learn salient features from molecular data. Recently, the field has seen a surge in the development of pretrained molecular embedding models. This trend mirrors the transformative impact of representation learning in other scientific domains, where models are pretrained on vast unlabeled datasets to produce powerful general-purpose embeddings. Notable successes include models based on self-supervised learning (SSL), such as DINO DINO and DINOv2 in computer vision, Sentence Transformers and TSDAE TSDAE for dense text representations in NLP, and multimodal architectures such as CLIP and BLIP-2 that connect images and text. Inspired by these advances, molecular models are similarly trained on massive chemical databases to generate universal embeddings.
@@ 1 Introduction
2 Molecular Embedding Approaches
@@ 2.1 Molecular fingerprints
Molecular fingerprints are feature extraction methods based on identifying small subgraphs within a molecule and detecting their presence or counting their occurrences, yielding binary and count variants, respectively. They can be broadly classified into substructural and hashed types scikit-fingerprints . Substructural fingerprints detect predefined patterns, such as functional groups or ring systems, that are typically determined by expert chemists. Hashed fingerprints, on the contrary, define general shapes of extracted subgraphs, convert them into numerical identifiers, and hash them using a modulo function into a fixed-length output vector. Common examples include circular neighborhoods in Extended Connectivity FingerPrint (ECFP) fp_ecfp , paths of length 4 in the Topological Torsion (TT) fp_topological_torsion , and the shortest paths between atom pairs in Atom Pair (AP) fingerprint.
@@
Although not task-adaptive, hashed fingerprints remain widely used in chemoinformatics and molecular machine learning due to their flexibility, computational efficiency, and consistently strong performance. In many cases, they continue to outperform more complex approaches, such as GNNs fingerprints_vs_gnns ; fingerprints_vs_gnns_2 ; MOLTOP ; peptides_fingerprints .
2.2 Graph neural networks
@@
Most Graph Neural Network (GNN) architectures follow a message-passing framework GNNs_message_passing . The initial embedding of each atom consists of elementary chemical descriptors such as the element type or charge. In each GNN layer, an atom receives embeddings from its neighbors (messages) and updates its own embedding accordingly, incorporating this information. Most molecular GNNs also include bond features and embeddings in this process. To obtain a whole-molecule embedding, atom embeddings are aggregated using a readout function, such as a channel-wise average or sum. The message-passing mechanism integrates both functional and structural information about the molecule in an end-to-end learning framework.
@@
Graph Isomorphism Network (GIN) GIN is one of the most widely used GNN architectures, as it was proven to be as expressive as the Weisfeiler-Lehman isomorphism test in distinguishing non-isomorphic graphs. GIN uses a two-layer multilayer perceptron (MLP) for atom embedding updates during message passing and employs sum readout. All GNN pretraining methods described below rely on a GIN backbone.
@@
Context Prediction (ContextPred) pretraining_GNNs_GIN was proposed as an SSL pretraining procedure for message-passing GNNs. For each atom, a $K$-hop neighborhood subgraph is defined and encoded as a vector by the GNN that is being pretrained. A corresponding context subgraph, located between $r_{1}$ and $r_{2}$ hops away, is encoded by a separate GNN. The pretraining objective is binary classification with negative sampling, where the model learns to distinguish correct (positive) neighborhood-context pairs from randomly sampled (negative) ones.
@@
GraphMVP GraphMVP combines contrastive and generative SSL, aligning molecular 2D and 3D representations, i.e., topological graphs and conformations. These two views are encoded using GIN and SchNet SchNet , respectively. In the contrastive setup, positive pairs consist of a molecule and its conformers, while the generative objective minimizes the variational autoencoder (VAE) loss between the two representations.
@@
GraphFP GraphFP introduces graph fragmentation for pretraining, employing both contrastive and predictive SSL. Frequent subgraph mining is used to decompose molecules into fragments. In contrastive learning, fragments and their constituent atoms form positive pairs, while atoms from unrelated fragments serve as negative pairs. The atom-encoding GNN is pretrained to generate atom-level embeddings (which can later be pooled), while a separate GNN encodes entire fragments. In addition, a predictive task is used to classify the presence of fragments, providing a multitask pretraining signal for the atom-encoding GNN.
@@
MolR MolR also uses contrastive learning, but incorporates chemical reaction information. Positive samples are constructed from known reactant-product pairs in reaction databases, while negative samples are created by randomly shuffling these pairs. All molecules are encoded using the same GNN being pretrained, based on the TAG architecture TAG .
@@
GEM GEM relies on predictive SSL, primarily using 3D molecular conformations. Pretraining tasks include predicting bond lengths, bond angles, and interatomic distances, with parts of the molecule masked as input. Additional auxiliary tasks involve predicting molecular fingerprint bits (MACCS and ECFP) based on the topological graph. Unlike the other architectures described above, which use only topological graphs at inference time, GEM requires 3D conformers as input, making both pretraining and inference computationally expensive due to the need for conformation generation.
2.3 Graph transformers
@@
Graph transformer architectures extend self‑attention mechanisms to molecular graphs, replacing or augmenting the localized message-passing of GNNs with global attention layers. Atoms are treated as tokens, and edges are encoded as pairwise biases injected into the attention scores, allowing information to propagate between any two atoms within a single layer. Compared to message-passing GNNs, this design more efficiently captures long-range dependencies and enables the incorporation of rich edge or distance features through additive or multiplicative attention biases.
@@
GROVER GROVER is a hybrid transformer-GNN model in which the attention heads are biased by chemoinformatics-derived edge features. It is pretrained using two self-supervised tasks: multioutput regression (MTR), predicting eight physicochemical descriptors on masked subgraphs, and multioutput classification (MTC), predicting the presence of functional groups (structural motifs). The graph-level embedding is obtained by concatenating a [CLS] token with pooled atomic embeddings.
@@
MAT MAT introduces distance-aware attention by incorporating adjacency and shortest-path kernels into the query-key dot product. Pretraining combines masked atom and bond prediction, analogous to masked language modeling (MLM) in NLP, with supervised regression on QM9 quantum-chemical properties, encouraging the encoder to learn both topological context and physicochemical properties.
@@
R-MAT RMAT builds on MAT by introducing relative positional encodings derived from graph distances and ring memberships, improving robustness to molecule size and graph sparsity. Inspired by GROVER, it uses subgraph masking as a pretraining objective, making it more challenging than single-atom masking of MAT, and resulting in superior performance.
@@
Uni-Mol UniMol and its extended version, Uni-Mol2 UniMol2 , apply graph transformers to 3D conformers using SE(3)-equivariant attention mechanisms. The encoder is jointly pretrained to reconstruct masked atoms, predict pairwise distances and angles, and distinguish positive from negative conformer–SMILES pairs. These objectives produce embeddings that integrate both molecular topology and 3D geometry.
2.4 Text transformers
@@
Molecules serialized as SMILES SMILES or SELFIES SELFIES strings can be processed directly using NLP-inspired architectures. Although pretraining can incorporate chemistry-specific biases, it is not strictly necessary, and the simplicity of NLP-derived pretraining is often a notable advantage of these models. In encoder–decoder architectures, only the encoder is used after pretraining to compute molecular embeddings.
@@
Continous Data-Driven Descriptors (CDDD) CDDD is an RNN-based architecture. It was trained using a purely text-based translation task, from random SMILES to canonical SMILES, calculated with RDKit RDKit .
@@
Chemformer Chemformer is based on BART encoder-decoder transformer BART , pretrained using denoising SSL objective. Given a SMILES string, it is first randomized (e.g., by starting at a different atom) and some tokens are randomly masked; the model is then trained to reconstruct the original SMILES.
@@
MolBERT MolBERT is an encoder-only transformer based on the original BERT architecture. It is pretrained using two objectives: masked language modeling (MLM), and multitask regression (MTR), in which the model predicts 200 physicochemical molecular properties computed with RDKit. This dual objective is intended to combine contextual learning from SMILES with molecular-level inductive biases.
@@
ChemBERTa ChemBERTa uses a RoBERTa-style encoder-only architecture. The authors explore two distinct pretraining variants: one based solely on MLM and another based solely on MTR. The MTR-only variant is reported to be simpler to train compared to MolBERT, and achieves better performance than both the MLM variant and MolBERT combined objective.
@@
MoLFormer MolFormer is another encoder-only SMILES transformer. Its authors aim to further simplify pretraining by relying entirely on a large pretraining corpus (derived from PubChem and ZINC) and standard NLP procedures, rather than domain-specific architectural changes. The model is based on BERT, with modifications including linear attention and rotary position embeddings, and is pretrained using MLM.
@@
SimSon SimSon uses an encoder-only backbone and is pretrained by contrastive SSL. Randomized SMILES strings of the same molecule form positive pairs, while strings from different molecules form negative pairs. Despite using such a simple pretraining procedure and a much smaller dataset, SimSon is shown to be competitive with the aforementioned architectures.
@@
ChemFM ChemFM is built on an encoder-only LLaMA LLaMa backbone and is pretrained using GPT-style causal language modeling (CLM). It is significantly larger than other models, with 1B and 3B parameter variants. The molecule embedding is derived from the output vector of the end-of-sequence ([EOS]) token.
@@
ChemGPT ChemGPT adopts a decoder-only architecture based on GPT-Neo GPTNeo , and is trained on SELFIES strings using CLM. The authors train multiple model variants up to 1.3B parameters, demonstrating that neural scaling laws hold for chemistry-focused text-based transformers as well.
@@
SELFormer SELFormer is based on an encoder-only RoBERTa architecture but uses SELFIES, rather than SMILES. Its authors assume that the more regular syntax of SELFIES facilitates learning and improves the capture of relevant molecular patterns. Pretraining uses the MLM objective.
2.5 Hybrid models
@@
Mol2Vec Mol2Vec combines graph-based ECFP4 descriptors with the text-based Word2Vec approach Word2Vec to learn context-aware token representations. It first computes integer ECFP identifiers for circular subgraphs centered on each atom, using radii 0 and 1. These subgraphs are treated as words in a sequence, with each one forming a token. A Word2Vec skip-gram model is then used to pretrain context-aware vector representations for each token. During inference, all subgraph vectors are summed to obtain a whole-molecule embedding.
@@
COATI COATI employs a multimodal encoder-decoder architecture, pretrained using CLIP-style CLIP multimodal contrastive learning. The 3D molecule conformer is encoded using an E(3)-equivariant GNN E3GNN , and its SMILES string with a RoFormer RoFormer . The model is pretrained with two objectives: contrastive learning between those two representations (treating embeddings of the same molecule as positive pairs), and autoregressive SMILES reconstruction using a decoder. At inference time, molecular embeddings are obtained from the SMILES encoder.
@@
CLAMP CLAMP is a multimodal encoder model trained with supervised contrastive learning based on ChEMBL bioactivity assay data. Each molecule is encoded via the concatenation of three molecular fingerprints (ECFP, RDKit, and MACCS) processed through a two-layer MLP, which serves as the final embedding. Assay descriptions are encoded using Latent Dirichlet Allocation (LDA). Positive pairs consist of a bioactive molecule and the description of the assay in which it is active, while negative pairs involve nonbioactive compounds. This design incorporates a bioactivity-specific inductive bias into the fingerprint-based representation.
3 Methods